OpenStudy (anonymous):

How do you solve 3x/5 - x = x/15 - 35/3 ?

ganeshie8 (ganeshie8):

multiply thru wid 15

OpenStudy (anonymous):

Multiply what?

ganeshie8 (ganeshie8):

\(\large \frac{3x}{5} - x = \frac{x}{15} - \frac{35}{3}\)

OpenStudy (anonymous):